# First-order definability of Campana Points and Darmon Points in algebraic function fields in one variable over number fields

### v1

@@ 4.1. Quadratic Pfister forms
$\begin{matrix}\langle\!\langle a_{1},\cdots,a_{n},b]]_{K}&:&V\times V&\to&K,\\ &&\left(v_{1},v_{2}\right)&\mapsto&q\left(v_{1}\right)-a_{n}q\left(v_{2}\right).\end{matrix}$ (1)
@@
Witt equivalence is stable under field extensions. This is, given a field $K$ and two quadratic forms $q:V\to K$ and $q^{\prime}:V^{\prime}\to K$, where $V$ and $V^{\prime}$ are finite-dimensional $K$-vector spaces such that $\left[q\right]=\left[q^{\prime}\right]\in WG\left(K\right)$, we then have $\left[q_{L}\right]=\left[q^{\prime}_{L}\right]\in WG\left(L\right)$, where for any quadratic form $\varphi:W\to K$ over $K$ where $W$ is a finite-dimensional $K$-vector space we define $\varphi_{L}:W\otimes_{K}L\to L$ to be the quadratic form over $L$ satisfying $\varphi\left(w\otimes\lambda\right)=\lambda^{2}\varphi\left(w\right)$ for every $\left(w,\lambda\right)\in W\times L$. Moreover, if $n\in\mathbb{Z}_{\geq 1}$ and $\left[\varphi\right]\in WG\left(K\right)$ equals the Witt class of some $n$-fold quadratic Pfister form over $K$, then $\left[\varphi_{L}\right]\in WG\left(L\right)$ equals the Witt class of some $n$-fold quadratic Pfister form over $L$. We then have the following well-defined group homomorphism:
@@ 4.1.1. Scharlau Transfer Maps
Assume $L$ is a finite field extension of a field $K$, and that $s:L\to K$ is any nonzero $K$-linear map. For any quadratic form $q$ over $L$, $s\circ q$ is a quadratic form over $K$. Moreover, if $q$ is non-singular, then so is $s\circ q$ ([12, Lemma 20.4(2)]), and if $q$ is hyperbolic, then so is $s\circ q$ ([12, Corollary 20.5(2)]). Therefore we have a well-defined group homomorphism $s_{\ast}:WG\left(L\right)\to WG\left(K\right)$ given by $\left[q\right]\mapsto\left[s\circ q\right]$ for every quadratic form $q$ over $L$.
@@
This is the same proof as in [4, Lemma 2.2]. The $K$-linear map $\mu:L\to L^{\ast}$ satisfying $\mu\left(a\right)\left(x\right)=s\left(ax\right)$ for every $a,x\in L$ is a monomorphism. Indeed, if $a\in\operatorname{ker}\left(\mu\right)$ then $s\left(ax\right)=0$ for every $x\in L$. If $a\neq 0$ then $s$ would be the zero element of $L^{\ast}$, a contradiction. Thus $a=0$ and $\mu$ is a monomorphism. Since $L$ is finite-dimensional over $K$ then $\dim_{K}\left(L\right)=\dim_{K}\left(L^{\ast}\right)$, thus $\mu$ is an isomorphism. We conclude that $t=\mu\left(a\right)$ for some $a\in L$; that is, $t\left(x\right)=s\left(ax\right)$ for all $x\in L$. We have $a\in L^{\times}$ because $t$ is not the zero map.
@@ 5. The $2$-dimension of a field
For the basic properties of the cohomological dimension we refer the reader to [32, §1.4]. In our case, we will be interested only in the following three properties:
@@ 6. Reciprocity theorem for quadratic Pfister forms
Follows directly from [4, Lemma 2.1]. $\blacksquare$
@@
$V_{q,v}$ $\coloneqq\left\{x\in V:q\left(v\left(x\right)\right)\geq 0\right\},$
$V_{q,v}^{0}$ $\coloneqq\left\{x\in V:q\left(v\left(x\right)\right)>0\right\}.$
@@
This is a particular case of [4, Proposition 2.3]. $\blacksquare$
@@
This is a particular case of [4, Lemma 3.8]. $\blacksquare$
@@
$WG^{d+1}\left(F\right)$${\bigoplus_{v\in\mathcal{V}\left(F/K\right)}WG^{d}\left(Fv\right)}$$WG^{d}\left(K\right)$}0{0} (2)
@@ 7. Describing our method and ensuring its efficacy
Our goal is to define Campana Points and Darmon Points over $F/K$, where $K$ is a number field and $F$ is an algebraic function field in one variable over $K$. Our method will heavily rely on the local-to-global properties of Pfister forms (see Section 6), for which we will require some hypotheses of regularity, linkage, and also the non-hyperbolicity of some Pfister forms. It is important to note that these hypotheses are not all met for every algebraic function field in one variable over a number field, therefore our first objective is to be able to reduce our problem to the case in which the hypotheses are actually met, the main ingredient for such a task being the following result proven in [4, Lemma 5.6]:
@@
Moreover, a weaker version of [4, Example 5.4(3)] can be stated as follows:
@@ 7.1. Coordinate polynomials and diophantine restrictions on finite extensions
For the first need; namely, the interpretation argument, we will use the theory of coordinate polynomials. For more on this we refer the reader to [36, §B.7].
@@
$\lambda=\sum_{j=1}^{n}\sigma_{j}\left(\lambda\right)\omega_{j}\text{ for every $\lambda\in L$}$ (3)
@@
$\sigma_{j}\left(P\left(\vec{x}\right)\right)=\sigma_{j}\left(P\right)\left(\vec{x}\right)\text{ for every $\vec{x}\in K^{m}$ and $j\in\left[1,n\right]\cap\mathbb{Z}$}.$ (4)
@@
Fix $m,r\in\mathbb{Z}_{\geq 1}$, $\vec{x}\in K^{m}$ and a polynomial $P\in L\left[X_{1},\cdots,X_{m},Y_{1},\cdots Y_{r}\right]$. For each $k\in\left[1,n\right]\mathbb{Z}$ define
@@
$P\left(\vec{x},\vec{y}\right)=P\left(\vec{x},\left(\sum_{j=1}^{n}\mathcal{Y}_{ij}\omega_{j}\right)_{i\in\left[1,r\right]\cap\mathbb{Z}}\right)\\ =\sum_{k=1}^{n}\sigma_{k}\left[P\left(\vec{x},\left(\sum_{j=1}^{n}\mathcal{Y}_{ij}\omega_{j}\right)_{i\in\left[1,r\right]\cap\mathbb{Z}}\right)\right]\omega_{k}=\sum_{k=1}^{n}P_{k}\left(\vec{x},\mathcal{Y}\right)\omega_{k},$ (5)
@@
Lemma 7.3 has the following immediate corollary, which allows us to restrict first-order definitions to base fields on finite extensions, and is a generalized version of [23, Lemma 6.3].
@@
Given $\vec{x}\in K^{m}$, we then have $\vec{x}\in D$ if and only if there exists $\vec{y}\in L^{r}$ such that $P\left(\vec{x},\vec{y}\right)=0$, which by Lemma 7.3 is equivalent to the existence of $\mathcal{Y}\in K^{r\times n}$ such that for every $k\in\left[1,n\right]\cap\mathbb{Z}$ we have $P_{k}\left(\vec{x},\mathcal{Y}\right)=0$. Since $K$ is not algebraically closed, [36, Lemma 1.2.3] says that there exists $Q\in K\left[X_{1},\cdots,X_{m},Y_{11},\cdots,Y_{rn}\right]$ which vanishes exactly at the common solutions of $\left\{P_{k}=0:k\in\left[1,n\right]\cap\mathbb{Z}\right\}$. We conclude that $D\cap K^{m}$ is definable over $K$ with the formula
@@ 7.2. Non-ramification of places by constant extensions
Since $L/K$ is finite and separable because $\operatorname{Char}\left(K\right)=0$, by the primitive element theorem ([20, Chapter V, Theorem 4.6]) there exists $\alpha\in L$ such that $L=K\left(\alpha\right)$, so that $FL=F\left(\alpha\right)$. Let $f\in K\left[x\right]\subseteq\mathcal{O}_{v}\left[x\right]$ be the minimal polynomial of $\alpha$ over $K$. Pick a monic irreducible factor $g$ of $f$ in $\mathcal{O}_{v}\left[x\right]$ having $\alpha$ as a root, so that $FL=F\left(\alpha\right)\cong F\left[x\right]/\left(g\right)$. Observe that the reduction of $g$ modulo $\mathfrak{m}_{v}$ is separable because $K$ is a characteristic-zero subfield of $Fv$, so Lemma 7.5 gives the desired conclusion. $\blacksquare$
@@ 8. First-order definitions
so that the image of $\beta$ under the second arrow of (2) equals $\left|S\right|\alpha$. Proposition 5.5 gives that $WG^{d+1}\left(K\right)$ is trivial, so by Lemma 4.8 the group $WG^{d+1}\left(K\right)$ is $2$-torsion, which together with the fact that $\left|S\right|$ is even gives that $\beta$ belongs to the kernel of the second arrow of (2). By the exactness of the sequence given by Theorem 6.5 there exists $\gamma\in WG^{d+1}\left(F\right)$ such that:
@@ 8.1. Some preliminary uniform first-order definitions
The assertion regarding the first set is a particular case of [2, Theorem 10.13], while the second assertion follows directly from the first one and the following equality, which is a direct consequence of [3, Lemma 5.4]:
@@
$\bigcap_{v\in S}\mathfrak{m}_{v}+\bigcap_{v\in T}\mathfrak{m}_{v}=\bigcap_{v\in S\cap T}\mathfrak{m}_{v},$ (6)
@@
Recalling that the formula $x\neq 0$ can be existentially defined as $\exists y\left(xy-1=0\right)$ and that any finite number of polynomial equations over $K$ can be reduced to a single one by [36, Lemma 1.2.3], then by Proposition 8.4 the formula
@@ 8.2. Campana points
$\forall\alpha_{1}\forall\alpha_{2}\forall\beta\forall\gamma\forall\alpha^{\prime}_{1}\forall\alpha^{\prime}_{2}\forall\beta^{\prime}\forall\gamma^{\prime}\\ \left[\begin{pmatrix}\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\alpha^{\prime}_{1}\alpha^{\prime}_{2}\gamma^{\prime}\left(1+4\beta^{\prime}\right)\neq 0\\ \Omega_{c,c^{\prime}}^{a_{1},a_{2},b,a_{1}^{\prime},a_{2}^{\prime},b^{\prime},F^{\prime}/L}\cap\Omega_{\gamma,\gamma^{\prime}}^{\alpha_{1},\alpha_{2},\beta,\alpha_{1}^{\prime},\alpha_{2}^{\prime},\beta^{\prime},F^{\prime}/L}=\emptyset\\ x\in\left(J_{\gamma,\gamma^{\prime}}\left(\alpha_{1},\alpha_{2},\beta,\alpha^{\prime}_{1},\alpha^{\prime}_{2},\beta^{\prime},F^{\prime}/L\right)\setminus\left\{0\right\}\right)^{-1}\end{pmatrix}\Rightarrow\right.\\ \left.x\in\left(J_{\gamma,\gamma^{\prime}}\left(\alpha_{1},\alpha_{2},\beta,\alpha^{\prime}_{1},\alpha^{\prime}_{2},\beta^{\prime},F^{\prime}/L\right)\setminus\left\{0\right\}\right)^{-n}\right].$ (7)
@@ 8.3. Darmon points
Second, given $m,n,r\in\mathbb{Z}_{\geq 1}$, polynomials $p\in F\left[x_{1},\cdots,x_{m},y_{1},\cdots,y_{n}\right]$ and $q\in F\left[z_{1},\cdots,z_{r}\right]$, consider the formula
$\left[\exists x_{1}\cdots\exists x_{m}\forall y_{1}\cdots\forall y_{n}\left(p\left(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n}\right)=0\right)\right]\vee\left[\exists z_{1}\cdots\exists z_{r}\left(q\left(z_{1},\cdots,z_{r}\right)=0\right)\right],$ (8)
which is the general shape of a disjunction between an $\exists\forall$-formula and an $\exists$-formula. Such a disjunction is always equivalent to an $\exists\forall$-formula. Indeed, adding another variable $u$, [36, Lemma 1.2.3] says that there exists $W\in F\left[x_{1},\cdots,x_{m},y_{1},\cdots,y_{n},z_{1},\cdots,z_{r},u\right]$ such that the system
$\begin{cases}p\left(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n}\right)&=0,\\ uq\left(z_{1},\cdots,z_{r}\right)-1&=0\end{cases}$
is equivalent to $W\left(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n},z_{1},\cdots,z_{r},u\right)=0$. Formula (8) is now equivalent to
$\exists x_{1}\cdots\exists x_{m}\exists z_{1}\cdots\exists z_{r}\forall y_{1}\cdots\forall y_{n}\forall u\left(W\left(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n},z_{1},\cdots,z_{r},u\right)\neq 0\right),$
@@
$\forall\alpha_{1}\forall\alpha_{2}\forall\beta\forall\gamma\left[\begin{pmatrix}\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\neq 0\\ \Omega_{c,c^{\prime}}^{a_{1},a_{2},b,a_{1}^{\prime},a_{2}^{\prime},b^{\prime},F^{\prime}/L}\cap\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\emptyset\\ \Delta_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\end{pmatrix}\Rightarrow\right.\\ \left.\exists y\exists z\left(y,z\in T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\wedge\operatorname{gcd}_{T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}}\left(y,z\right)=1\wedge y=xz^{n}\right)\right].$ (9)
@@
$\exists y\exists z\left(y,z\in T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\wedge\operatorname{gcd}_{T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}}\left(y,z\right)=1\wedge y=xz^{n}\right)$ (10)
@@
$\begin{pmatrix}\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\neq 0\\ \Omega_{c,c^{\prime}}^{a_{1},a_{2},b,a_{1}^{\prime},a_{2}^{\prime},b^{\prime},F^{\prime}/L}\cap\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\emptyset\\ \Delta_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\end{pmatrix}$ (11)
@@
Given $x\in F^{\prime}$ such that (9) holds on $x$ and $w\in\mathcal{V}\left(F^{\prime}/L\right)\setminus S^{\prime}$ such that $w\left(x\right)\not\in\mathbb{Z}_{\geq 0}$, we want to show that $n\mid w\left(x\right)$. Fixing $w^{\prime}\in\mathcal{V}\left(F^{\prime}/L\right)\setminus\left(S^{\prime}\cup\left\{w\right\}\right)$, by Proposition 8.3 there exist $\alpha_{1},\alpha_{2},\beta,\gamma\in F$ with $\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\neq 0$ and $\Delta_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\left\{w,w^{\prime}\right\}$. Now formula (9) implies that $y=xz^{n}$ for some relatively prime $x,y\in\mathcal{O}_{w}\cap\mathcal{O}_{w^{\prime}}$. Applying $w$ to this equality we get $w\left(y\right)\equiv w\left(x\right)\pmod{n}$, so it suffices to show that $w\left(y\right)=0$. Since $x,y\in\mathcal{O}_{w}\cap\mathcal{O}_{w^{\prime}}$ are relatively prime, it suffices to show that $w\left(z\right)>0$. Indeed, if $w\left(z\right)=0$ then $w\left(y\right)=w\left(x\right)$, which is impossible because $w\left(x\right)<0\leq w\left(y\right)$.
@@
Since $x\in D_{F^{\prime}/L,S^{\prime},n}$, we get $w\left(x\right)\in\mathbb{Z}_{\geq 0}\cup n\mathbb{Z}$ for all $w\in\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}$.

### v2

@@ 4.1. Quadratic Pfister forms
(1) $\begin{matrix}\langle\!\langle a_{1},\cdots,a_{n},b]]_{K}&:&V\times V&\to&K,\\ &&\left(v_{1},v_{2}\right)&\mapsto&q\left(v_{1}\right)-a_{n}q\left(v_{2}\right).\end{matrix}$
@@
Witt equivalence is stable under field extensions. This is, given a field $K$ and two quadratic forms $q:V\to K$ and $q^{\prime}:V^{\prime}\to K$, where $V$ and $V^{\prime}$ are finite-dimensional $K$-vector spaces such that $\left[q\right]=\left[q^{\prime}\right]\in WG\left(K\right)$, we then have $\left[q_{L}\right]=\left[q^{\prime}_{L}\right]\in WG\left(L\right)$, where for any quadratic form $\varphi:W\to K$ over $K$ where $W$ is a finite-dimensional $K$-vector space we define $\varphi_{L}:W\otimes_{K}L\to L$ to be the quadratic form over $L$ satisfying $\varphi_{L}\left(w\otimes\lambda\right)=\lambda^{2}\varphi\left(w\right)$ for every $\left(w,\lambda\right)\in W\times L$. Moreover, if $n\in\mathbb{Z}_{\geq 1}$ and $\left[\varphi\right]\in WG\left(K\right)$ equals the Witt class of some $n$-fold quadratic Pfister form over $K$, then $\left[\varphi_{L}\right]\in WG\left(L\right)$ equals the Witt class of some $n$-fold quadratic Pfister form over $L$. We then have the following well-defined group homomorphism:
@@ 4.1.1. Scharlau Transfer Maps
Assume $L$ is a finite field extension of a field $K$, and that $s:L\to K$ is any nonzero $K$-linear map. For any quadratic form $q$ over $L$, $s\circ q$ is a quadratic form over $K$. Moreover, if $q$ is non-singular, then so is $s\circ q$ ([undefk, Lemma 20.4(2)]), and if $q$ is hyperbolic, then so is $s\circ q$ ([undefk, Corollary 20.5(2)]). Therefore we have a well-defined group homomorphism $s_{\ast}:WG\left(L\right)\to WG\left(K\right)$ given by $\left[q\right]\mapsto\left[s\circ q\right]$ for every quadratic form $q$ over $L$.
@@
This is the same proof as in [undefc, Lemma 2.2]. The $K$-linear map $\mu:L\to L^{\ast}$ satisfying $\mu\left(a\right)\left(x\right)=s\left(ax\right)$ for every $a,x\in L$ is a monomorphism. Indeed, if $a\in\operatorname{ker}\left(\mu\right)$ then $s\left(ax\right)=0$ for every $x\in L$. If $a\neq 0$ then $s$ would be the zero element of $L^{\ast}$, a contradiction. Thus $a=0$ and $\mu$ is a monomorphism. Since $L$ is finite-dimensional over $K$ then $\dim_{K}\left(L\right)=\dim_{K}\left(L^{\ast}\right)$, thus $\mu$ is an isomorphism. We conclude that $t=\mu\left(a\right)$ for some $a\in L$; that is, $t\left(x\right)=s\left(ax\right)$ for all $x\in L$. We have $a\in L^{\times}$ because $t$ is not the zero map.
@@ 5. The $2$-dimension of a field
For the basic properties of the cohomological dimension we refer the reader to [undefae, §1.4]. In our case, we will be interested only in the following three properties:
@@ 6. Reciprocity theorem for quadratic Pfister forms
Follows directly from [undefc, Lemma 2.1]. $\blacksquare$
@@
$V_{q,v}$ $\coloneqq\left\{x\in V:v\left(q\left(x\right)\right)\geq 0\right\},$
$V_{q,v}^{0}$ $\coloneqq\left\{x\in V:v\left(q\left(x\right)\right)>0\right\}.$
@@
This is a particular case of [undefc, Proposition 2.3]. $\blacksquare$
@@
This is a particular case of [undefc, Lemma 3.8]. $\blacksquare$
@@
(2) $WG^{d+1}\left(F\right)$${\bigoplus_{v\in\mathcal{V}\left(F/K\right)}WG^{d}\left(Fv\right)}$$WG^{d}\left(K\right)$}0{0}
@@ 7. Describing our method and ensuring its efficacy
Our goal is to define Campana Points and Darmon Points over $F/K$, where $K$ is a number field and $F$ is an algebraic function field in one variable over $K$. Our method will heavily rely on the local-to-global properties of Pfister forms (see Section 6), for which we will require some hypotheses of regularity, linkage, and also the non-hyperbolicity of some Pfister forms. It is important to note that these hypotheses are not all met for every algebraic function field in one variable over a number field, therefore our first objective is to be able to reduce our problem to the case in which the hypotheses are actually met, the main ingredient for such a task being the following result proven in [undefc, Lemma 5.6]:
@@
Moreover, a weaker version of [undefc, Example 5.4(3)] can be stated as follows:
@@ 7.1. Coordinate polynomials and diophantine restrictions on finite extensions
For the first need; namely, the interpretation argument, we will use the theory of coordinate polynomials. For more on this we refer the reader to [undefai, §B.7].
@@
(3) $\lambda=\sum_{j=1}^{n}\sigma_{j}\left(\lambda\right)\omega_{j}\text{ for every $\lambda\in L$}$
@@
(4) $\sigma_{j}\left(P\left(\vec{x}\right)\right)=\sigma_{j}\left(P\right)\left(\vec{x}\right)\text{ for every $\vec{x}\in K^{m}$ and $j\in\left[1,n\right]\cap\mathbb{Z}$}.$
@@
Fix $m,r\in\mathbb{Z}_{\geq 1}$, $\vec{x}\in K^{m}$ and a polynomial $P\in L\left[X_{1},\cdots,X_{m},Y_{1},\cdots Y_{r}\right]$. For each $k\in\left[1,n\right]\cap\mathbb{Z}$ define
@@
(5) $P\left(\vec{x},\vec{y}\right)=P\left(\vec{x},\left(\sum_{j=1}^{n}\mathcal{Y}_{ij}\omega_{j}\right)_{i\in\left[1,r\right]\cap\mathbb{Z}}\right)\\ =\sum_{k=1}^{n}\sigma_{k}\left[P\left(\vec{x},\left(\sum_{j=1}^{n}\mathcal{Y}_{ij}\omega_{j}\right)_{i\in\left[1,r\right]\cap\mathbb{Z}}\right)\right]\omega_{k}=\sum_{k=1}^{n}P_{k}\left(\vec{x},\mathcal{Y}\right)\omega_{k},$
@@
Lemma 7.3 has the following immediate corollary, which allows us to restrict first-order definitions to base fields on finite extensions, and is a generalized version of [undefv, Lemma 6.3].
@@
Given $\vec{x}\in K^{m}$, we then have $\vec{x}\in D$ if and only if there exists $\vec{y}\in L^{r}$ such that $P\left(\vec{x},\vec{y}\right)=0$, which by Lemma 7.3 is equivalent to the existence of $\mathcal{Y}\in K^{r\times n}$ such that for every $k\in\left[1,n\right]\cap\mathbb{Z}$ we have $P_{k}\left(\vec{x},\mathcal{Y}\right)=0$. Since $K$ is not algebraically closed, [undefai, Lemma 1.2.3] says that there exists $Q\in K\left[X_{1},\cdots,X_{m},Y_{11},\cdots,Y_{rn}\right]$ which vanishes exactly at the common solutions of $\left\{P_{k}=0:k\in\left[1,n\right]\cap\mathbb{Z}\right\}$. We conclude that $D\cap K^{m}$ is definable over $K$ with the formula
@@ 7.2. Non-ramification of places by constant extensions
Since $L/K$ is finite and separable because $\operatorname{Char}\left(K\right)=0$, by the primitive element theorem ([undefs, Chapter V, Theorem 4.6]) there exists $\alpha\in L$ such that $L=K\left(\alpha\right)$, so that $FL=F\left(\alpha\right)$. Let $f\in K\left[x\right]\subseteq\mathcal{O}_{v}\left[x\right]$ be the minimal polynomial of $\alpha$ over $K$. Pick a monic irreducible factor $g$ of $f$ in $\mathcal{O}_{v}\left[x\right]$ having $\alpha$ as a root, so that $FL=F\left(\alpha\right)\cong F\left[x\right]/\left(g\right)$. Observe that the reduction of $g$ modulo $\mathfrak{m}_{v}$ is separable because $K$ is a characteristic-zero subfield of $Fv$, so Lemma 7.5 gives the desired conclusion. $\blacksquare$
@@ 8. First-order definitions
so that the image of $\beta$ under the second arrow of (2) equals $\left|S\right|\alpha$. Proposition 5.5 gives that $WG^{d+1}\left(K\right)$ is trivial, so by Lemma 4.8 the group $WG^{d}\left(K\right)$ is $2$-torsion, which together with the fact that $\left|S\right|$ is even gives that $\beta$ belongs to the kernel of the second arrow of (2). By the exactness of the sequence given by Theorem 6.5 there exists $\gamma\in WG^{d+1}\left(F\right)$ such that:
@@ 8.1. Some preliminary uniform first-order definitions
The assertion regarding the first set is a particular case of [undefa, Theorem 10.13], while the second assertion follows directly from the first one and the following equality, which is a direct consequence of [undefb, Lemma 5.4]:
@@
(6) $\bigcap_{v\in S}\mathfrak{m}_{v}+\bigcap_{v\in T}\mathfrak{m}_{v}=\bigcap_{v\in S\cap T}\mathfrak{m}_{v},$
@@
Recalling that the formula $x\neq 0$ can be existentially defined as $\exists y\left(xy-1=0\right)$ and that any finite number of polynomial equations over $K$ can be reduced to a single one by [undefai, Lemma 1.2.3], then by Proposition 8.4 the formula
@@ 8.2. Campana points
(7) $\forall\alpha_{1}\forall\alpha_{2}\forall\beta\forall\gamma\forall\alpha^{\prime}_{1}\forall\alpha^{\prime}_{2}\forall\beta^{\prime}\forall\gamma^{\prime}\\ \left[\begin{pmatrix}\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\alpha^{\prime}_{1}\alpha^{\prime}_{2}\gamma^{\prime}\left(1+4\beta^{\prime}\right)\neq 0\\ \Omega_{c,c^{\prime}}^{a_{1},a_{2},b,a_{1}^{\prime},a_{2}^{\prime},b^{\prime},F^{\prime}/L}\cap\Omega_{\gamma,\gamma^{\prime}}^{\alpha_{1},\alpha_{2},\beta,\alpha_{1}^{\prime},\alpha_{2}^{\prime},\beta^{\prime},F^{\prime}/L}=\emptyset\\ x\in\left(J_{\gamma,\gamma^{\prime}}\left(\alpha_{1},\alpha_{2},\beta,\alpha^{\prime}_{1},\alpha^{\prime}_{2},\beta^{\prime},F^{\prime}/L\right)\setminus\left\{0\right\}\right)^{-1}\end{pmatrix}\Rightarrow\right.\\ \left.x\in\left(J_{\gamma,\gamma^{\prime}}\left(\alpha_{1},\alpha_{2},\beta,\alpha^{\prime}_{1},\alpha^{\prime}_{2},\beta^{\prime},F^{\prime}/L\right)\setminus\left\{0\right\}\right)^{-n}\right].$
@@ 8.3. Darmon points
Second, given $m,n,r,k,\ell\in\mathbb{Z}_{\geq 0}$, polynomials $p\in F\left[x_{1},\cdots,x_{m},y_{1},\cdots,y_{n},u_{1},\cdots,u_{k}\right]$ and $q\in F\left[z_{1},\cdots,z_{r},v_{1},\cdots,v_{\ell}\right]$, consider the formula
(8) $\left[\exists x_{1}\cdots\exists x_{m}\forall y_{1}\cdots\forall y_{n}\left(p\left(\vec{x},\vec{y},\vec{u}\right)=0\right)\right]\vee\left[\exists z_{1}\cdots\exists z_{r}\left(q\left(\vec{z},\vec{v}\right)\neq 0\right)\right],$
which is the general shape of a disjunction between an $\exists\forall$-formula and an $\exists$-formula. Such a disjunction is always equivalent to an $\exists\forall$-formula. Indeed, adding another variable $u$, [undefai, Lemma 1.2.3] says that there exists $W\in F\left[\vec{x},\vec{y},\vec{z},\vec{u},\vec{v},u\right]$ such that the system
$\begin{cases}p\left(\vec{x},\vec{y},\vec{u}\right)&=0,\\ uq\left(\vec{z},\vec{v}\right)-1&=0\end{cases}$
is equivalent to $W\left(\vec{x},\vec{y},\vec{z},\vec{u},\vec{v},u\right)=0$. Formula (8) is now equivalent to
$\exists x_{1}\cdots\exists x_{m}\exists z_{1}\cdots\exists z_{r}\forall y_{1}\cdots\forall y_{n}\forall u\left(W\left(\vec{x},\vec{y},\vec{z},\vec{u},\vec{v},u\right)\neq 0\right),$
@@
(9) $\forall\alpha_{1}\forall\alpha_{2}\forall\beta\forall\gamma\left[\begin{pmatrix}\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\neq 0\\ \Omega_{c,c^{\prime}}^{a_{1},a_{2},b,a_{1}^{\prime},a_{2}^{\prime},b^{\prime},F^{\prime}/L}\cap\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\emptyset\\ \Delta_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\end{pmatrix}\Rightarrow\right.\\ \left.\exists y\exists z\left(y,z\in T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\wedge\operatorname{gcd}_{T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}}\left(y,z\right)=1\wedge y=xz^{n}\right)\right].$
@@
(10) $\exists y\exists z\left(y,z\in T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\wedge\operatorname{gcd}_{T_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}}\left(y,z\right)=1\wedge y=xz^{n}\right)$
@@
(11) $\begin{pmatrix}\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\neq 0\\ \Omega_{c,c^{\prime}}^{a_{1},a_{2},b,a_{1}^{\prime},a_{2}^{\prime},b^{\prime},F^{\prime}/L}\cap\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\emptyset\\ \Delta_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}\end{pmatrix}$
@@
Given $x\in F^{\prime}$ such that (9) holds on $x$ and $w\in\mathcal{V}\left(F^{\prime}/L\right)\setminus S^{\prime}$ such that $w\left(x\right)\not\in\mathbb{Z}_{\geq 0}$, we want to show that $n\mid w\left(x\right)$. Fixing $w^{\prime}\in\mathcal{V}\left(F^{\prime}/L\right)\setminus\left(S^{\prime}\cup\left\{w\right\}\right)$, by Proposition 8.3 there exist $\alpha_{1},\alpha_{2},\beta,\gamma\in F$ with $\alpha_{1}\alpha_{2}\gamma\left(1+4\beta\right)\neq 0$ and $\Delta_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\left\{w,w^{\prime}\right\}$. Now formula (9) implies that $y=xz^{n}$ for some relatively prime $y,z\in\mathcal{O}_{w}\cap\mathcal{O}_{w^{\prime}}$. Applying $w$ to this equality we get $w\left(y\right)\equiv w\left(x\right)\pmod{n}$, so it suffices to show that $w\left(y\right)=0$. Since $y,z\in\mathcal{O}_{w}\cap\mathcal{O}_{w^{\prime}}$ are relatively prime, it suffices to show that $w\left(z\right)>0$. Indeed, if $w\left(z\right)=0$ then $w\left(y\right)=w\left(x\right)$, which is impossible because $w\left(x\right)<0\leq w\left(y\right)$.
@@
Since $x\in D_{F^{\prime}/L,S^{\prime},n}$, we get $w\left(x\right)\in\mathbb{Z}_{\geq 0}\cup n\mathbb{Z}$ for all $w\in\Delta_{\gamma}^{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}=\Delta_{\alpha_{1},\alpha_{2},\beta,F^{\prime}/L}$.